# MVGT: A Multi-view Graph Transformer Based on Spatial Relations for EEG Emotion Recognition

### v2

@@ Abstract
Electroencephalography (EEG), a medical imaging technique that captures scalp electrical activity of brain structures via electrodes, has been widely used in affective computing. The spatial domain of EEG is rich in affective information. However, few of the existing studies have simultaneously analyzed EEG signals from multiple perspectives of geometric and anatomical structures in spatial domain. In this paper, we propose a multi-view Graph Transformer (MVGT) based on spatial relations, which integrates information from the temporal, frequency and spatial domains, including geometric and anatomical structures, so as to enhance the expressive power of the model comprehensively. We incorporate the spatial information of EEG channels into the model as encoding, thereby improving its ability to perceive the spatial structure of the channels. Meanwhile, experimental results based on publicly available datasets demonstrate that our proposed model outperforms state-of-the-art methods in recent years. In addition, the results also show that the MVGT could extract information from multiple domains and capture inter-channel relationships in EEG emotion recognition tasks effectively.
@@ I Introduction
Affective computing is commonly employed for the analysis of emotional states through Human-Computer Interaction (HCI) systems, which collect multimodal data from subjects, including voice signal, self-report, body gesture and physiological signals. Compared to other modalities, physiological signals have certain advantages. These signals are directly captured from the subjects’ mental states, thus prevent subjects from disguising or hiding. The physiological signals commonly used to measure emotions are electroencephalography (EEG), electrocardiography (ECG), electromyography (EMG), and galvanic skin response (GSR), etc., among which EEG is often utilized for analyzing cognitive functions of human brain. Electrical signals from brain neurons are collected using the EEG method, which involves placing dry and noninvasive electrodes on the scalp [1]. Nowadays, due to its high temporal resolution, portability, and affordability, this method is widely employed to study brain changes in response to emotional stimuli [2].
@@
Traditional EEG features are mainly divided into three kinds, i.e., time domain, frequency domain, and time-frequency domain features. Given the signal-to-noise ratio and substantial fluctuations inherent in EEG signals, frequency domain features are commonly used for EEG-based emotion recognition tasks. The typical approach involves decomposing the raw signals into five frequency bands: $\delta$, $\theta$, $\alpha$, $\beta$, $\gamma$. Frequency domain features, such as power spectral density (PSD) [3], differential entropy (DE) [4, 5], differential asymmetry (DASM) [6] and rational asymmetry (RASM) [7], are subsequently extracted from each frequency band respectively.
@@
The spatial structure of the brain also contains rich emotional information. Emotional states may involve distributed circuits rather than considering a single brain region in isolation [8]. Asymmetry between the left and right hemispheres can reflect changes in valence and arousal [9]. Recent studies have highlighted the importance of utilizing spatial domain information. Li et al. [10] introduced recurrent neural networks to learn the asymmetric differences between the left and right hemispheres. Li et al. [11] also utilized hierarchical neural networks to learn both regional and global information of spatial-temporal EEG features. Graph Neural Networks (GNN) are emerging as a powerful tool for analyzing spatial information in EEG emotion recognition. Song et al. [12] dynamically learned relationships between EEG channels using graph convolutional networks (GCN). Zhong et al. [13] incorporated asymmetry of the hemispheres into the adjacency matrix to model graph structure. Li et al. [14] also utilized adaptive GCNs that integrate multi-domain information to learn relationships between channels. Ding et al. [15] incorporated lobe information as prior knowledge into the GNN. Jiang et al. [16] proposed an elastic graph Transformer to extract emotional information. Although these methods have achieved excellent performance in emotion recognition, they have a common issue: they all rely on GNNs based on neighborhood aggregation schemes which may pose potential risks such as over-smoothing [17, 18, 19], under-reaching [20], and over-squashing [21]. Additionally, these methods do not comprehensively consider the geometric and anatomical structure information of the brain.
The main contributions of this paper lie in three aspects:
  • $\bullet$
    We propose a multi-view graph transformer based on spatial relations (MVGT), fusing information from multiple perspectives including temporal, frequency, and spatial domains.
  • $\bullet$
@@
    Our method, based on Graph Transformer, mitigates the potential risks of over-smoothing, under-reaching and over-squashing occurring in traditional GNNs. Additionally, it enhances the model’s expressive power in emotion recognition by introducing spatial structural encoding based on geometric and brain lobe information.
  • $\bullet$
    Extensive experiments conducted on public datasets SEED and SEED-IV show the effectiveness of our model in emotion classification tasks.
@@ II Related Work
In this section, we review the related work from the perspectives of EEG-based emotion recognition and graph transformer.
@@ II-A EEG-based Emotion Recognition
Along with the development of deep learning, increasingly complex models with rich expressive abilities have emerged and have been extensively utilized in EEG signal analysis. Zheng et al. [5] employed deep belief networks to analyze important frequency domain components and effective channels based on the learned parameters. Song et al. [12] used a GCN method based on Chebyshev polynomials [24] to dynamically learn the representations of EEG signals. Zhong et al. [13] innovatively incorporated the asymmetric information of the hemispheres as prior knowledge into the adjacency matrix in 3D space. The reasonable combination of the multi-domain information contributes to improving the accuracy in the emotion recognition task. Li et al. [14] proposed an adaptive GCN that integrates the temporal domain, frequency domain, and functional connectivity. Ding et al. [15], inspired by neuroscience research, combined intra-region convolution and inter-region convolution based on brain lobe regions to learn cognitive patterns. Jiang et al. [16] utilized the advantages of GCN in the spatial domain and Transformer in the temporal domain to improve the accuracy of emotion classification.
@@ II-B Graph Transformer
The GNNs used in the above methods are based on neighborhood aggregation schemes. However, classical GNNs based on message passing (MPGNNs) may lead to over-smoothing [17, 18, 19], under-reaching [20], and may also fail to fit long-range signals due to over-squashing [21], which limit the expressive power of the model. Graph transformers (GTs) alleviate such effects as they have a global receptive field [25]. Nevertheless, without sufficiently expressive structural and positional encodings, GTs cannot capture effective graph structures [26]. Dwivedi et al. [27] utilized eigenvectors of graph Laplacian as position encodings in fully connected Graph Transformers and integrated edge features into the attention mechanism. Building on this, SAN [28] used a full Laplacian spectrum to learn the positional encodings for each node. Graphormer [29, 30] employed node centrality and node distance metric to implement structural and relative positional encodings, achieving state-of-the-art performance on challenging graph datasets. In EEG emotion recognition, Li et al. [31] innovatively combined a masked autoencoder based on self-supervised learning with a CNN-Transformer hybrid structure, effectively improving classification accuracy. However, this method only used the sine-cosine positional encoding, limiting the Transformer’s ability to learn spatial information.
@@ III-A Graph Neural Network (GNN)
Let $G=\left(V,E\right)$ define a graph, where $V=\left\{v_{1},v_{2},\cdots,v_{n}\right\}$ represents the nodes in the graph, and $E=\left\{e_{1},e_{2},\cdots,e_{m}\right\}$ is the edges between the nodes. The representation of node $v_{i}$ is denoted as $x_{i}\in\mathbb{R}^{d}$. Most existing GNNs [32, 17, 33, 34, 35] adopt neighborhood aggregation schemes, iteratively aggregating representations of its first or higher-order neighbors, followed by using backpropagation (BP) to learn task-driven feature encodings. We define the representation of node $v_{i}$ at the $l$ iteration as $h_{i}^{l}$ and define $h_{i}^{0}=x_{i}$. The $l$-th iteration can be represented as:
$a_{i}^{l}$ $=\text{AGGREGATE}^{l}\left(\left\{\varphi_{\theta}(h_{j}^{l},e_{ji}^{l}):j\in\mathcal{N}(v_{i})\right\}\right)$ (1)
$h_{i}^{l}$ $=\text{UPDATE}^{l}\left(h_{i}^{l-1},a_{i}^{l}\right)$ (2)
where $\varphi_{\theta}$ represents a differentiable function used for feature transformation of node and edge information. $\mathcal{N}(v_{i})$ is the set of neighbors of $v_{i}$. The AGGREGATE function is used to aggregate the transformed representation using a differentiable, permutation invariant function, (such as mean, sum, max, etc.). The goal of UPDATE function is to integrate the information from neighbors into the node representation. For graph classification, the READOUT operation is typically used to obtain a representation of the entire graph, which is then fed into a classifier to determine the graph label.
@@ III-B Graphormer
The Transformer [36] is undeniably one of the most popular deep neural network architectures today, driving significant advancements in natural language processing and computer vision. From the perspective of GNNs, Transformer can be interpreted as a GNN acting on a fully connected graph. Therefore, it is reasonable and feasible to use Transformer to address tasks on graph data. The ability to properly incorporate the structural information of graphs into the model is the key for leveraging its expressive power. Graphormer [29, 30] can go beyond classical MPGNNs in expressive power and achieves state-of-the-art performance on large molecular benchmarks. Graphormer incorporates centrality encoding into the graph data and integrates spatial encoding, edge encoding into the attention mechanism, which can be expressed as:
$A_{ij}=\frac{\left(h_{i}W_{Q}\right)\left(h_{j}W_{K}\right)^{T}}{\sqrt{d}}+b_{\phi(v_{i},v_{j})}+c_{ij},$ (3)
@@ IV METHODS
In this section, we introduce the methods employed in the EEG emotion recognition task. Firstly, we elaborate on the embedding for temporal information. Secondly, leveraging the spatial geometry and physiological anatomy of the brain, we propose two novel and simple designs of encoding that enable the model to adaptively learn the inter-channel correlations. Finally, we present the detailed implementations of MVGT.
@@ IV-A Problem Definition
EEG signals can be represented as a two-dimensional matrix with respect to channels and time. Given that channels exhibit spatial structure, they can be structured into fully connected graph data $G=\left(V,E\right)$, where $V$ denotes the nodes in the graph, representing EEG channels, and $E$ denotes the edges, representing the connections between channels. The features of the nodes at time $t$ are denoted by $\bm{X}_{t}=\left(\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}\right)\in\mathbb{R}^{n\times d}$, where $n=|V|$ represents the number of nodes and $d$ represents the feature dimension.
@@ IV-B Temporal Embedding
EEG signals have high temporal resolution and contain rich temporal information. Because of the multi-electrode acquisition method, EEG signals can be regarded as multivariate time series. When processing time series, the embedding of temporal information are crucial. EmoGT [16] treats the features of different channels at the same time points as tokens and employs an attention mechanism to extract temporal correlations between them. Due to the different anisotropic volume conduction characteristics [37] in human brain tissues, there may be temporal delays between different channels, which in turn leads to time-unaligned events at a single moment thus causing performance degradation. MD-AGCN [14] utilizes convolutional neural networks to extract temporal information along the time axis from continuous EEG segments, with the receptive field limited by the size of the convolution kernel. Inspired by iTransformer [38], we broaden the receptive field by considering the entire time series as an embedded token rather than a single time point. First, following the methods of MD-AGCN and EmoGT, we use overlapping sliding windows of size $T$ to segment EEG signals along the time axis and use these segments as tokens, which are then fed into the attention module in the form of continuous segments. After processing with sliding windows, we obtain $\tilde{\bm{X}}=\left(\tilde{\bm{X}}_{1},\tilde{\bm{X}}_{2},\cdots,\tilde{\bm{X}}_{S}\right)\in\mathbb{R}^{S\times n\times Tf}$, where $S$ denotes the number of continuous EEG segments, $n$ is the number of channels, and $f$ denotes the dimension of frequency domain features.
@@ IV-C Spatial Encoding
The special structure of the brain encompasses rich spatial information. Fully exploiting structural information is beneficial for the recognition and analysis of cognitive patterns in the brain. Therefore, to better identify emotional patterns in emotion classification tasks, we employed two simple but effective methods of spatial encoding: brain region encoding and geometric structure encoding.
@@ IV-C1 Brain Region Encoding
Neuroscience research demonstrated that the activation of a specific brain region often leads to the concurrent activation of related brain regions responsible for the same high-level cognition [40]. In EEG emotion recognition, incorporating relevant neuroscience findings can typically enhance recognition accuracy. RGNN [13] integrates the asymmetry of neural activity between the left and right hemispheres as prior knowledge into the adjacency matrix, effectively enhancing recognition accuracy. BiHDM [10] improves emotion pattern recognition performance by learning the differences between the left and right hemispheres. LGGNet [15] divides EEG channels into different regions and combines local intra-region convolution with global inter-region convolution, achieving good results on the DEAP [41] dataset. With reference to the three divisions of LGGNet, we adopt four brain region divisions, which divide the EEG channels into different regions based on a prior knowledge, aiming to incorporate the brain region information into the model. The division schemes consist of: (1) We divide the regions based on the anatomical structure of the brain and implement LOBE scheme. (2) To further investigate the expressive power of brain region encoding, we conduct a detailed division of brain lobes according to the 10-20 system based on electrode positions, employing the GENERAL scheme. (3) Asymmetric EEG activity in the frontal lobe can be utilized for discriminating valence changes [9]. The left frontal lobe exhibits a stronger correlation with joy and happy, while the right frontal lobe is more strongly correlated with fear and sadness. Thus we further divide the frontal lobe region into two symmetrical regions to obtain the FRONTAL scheme. (4) According to the symmetry of brain structure [42], we make a finer division of the brain lobe regions, defining the HEMISPHERE scheme.
The four schemes mentioned above are showed in Fig. 2. In terms of specific implementation, we assign each electrode a brain region tag, then project the tags into an embedding space using a learnable projection function, and simply add the embeddings to the node features. The encoding of node $i$ is represented as follows:
@@
where $\bm{W}_{\mathcal{X}}\in\mathbb{R}^{Tf\times d}$ is a learnable projection function, and $d$ represents the dimension of the embedding. Through the above encoding method, we integrate the information of the brain’s physiological anatomy into the model.

### v3

@@ Abstract
Electroencephalography (EEG), a medical imaging technique that captures scalp electrical activity of brain structures via electrodes, has been widely used in affective computing. The spatial domain of EEG data is rich in affective information. However, few of the existing studies have simultaneously analyzed EEG signals from the perspectives of geometric and anatomical structures in spatial domain. In this paper, we propose a multi-view graph transformer (MVGT) based on spatial relations, which integrates information from the temporal, frequency and spatial domains, including geometric and anatomical structures, so as to enhance the expressive power of the model comprehensively. We encode the spatial information of EEG channels into the model, thereby improving its ability to comprehend the spatial structure of the channels. Experimental results from publicly available datasets demonstrate that our proposed model outperforms state-of-the-art baseline methods. Furthermore, the results also show that the MVGT could extract information from multiple domains and capture inter-channel relationships in EEG-based emotion recognition tasks effectively.
@@ I Introduction
Affective computing is commonly employed for the analysis of emotional states through Human-Computer Interaction (HCI) systems, which collect multimodal data from subjects, including voice signals, self-reports, body gestures and physiological signals. Compared to other modalities, physiological signals have certain advantages. These signals are directly captured from the subjects’ mental states, thus prevent subjects from disguising or hiding. The physiological signals commonly used to measure emotions are electroencephalography (EEG), electrocardiography (ECG), electromyography (EMG), and galvanic skin response (GSR), etc., among which EEG is often utilized to analyze the cognitive functions of human brain. Electrical signals from brain neurons are collected through dry and noninvasive electrodes placed on the scalp [1]. Nowadays, due to its high temporal resolution, portability, and affordability, this method is widely employed to study brain changes in response to emotional stimuli [2].
@@
Traditional EEG features are mainly divided into three categories, i.e., time domain, frequency domain, and time-frequency domain features. Given the low signal-to-noise ratio and substantial fluctuations inherent in EEG signals, frequency domain features are often used for EEG-based emotion recognition tasks. The typical approach involves decomposing the raw signals into five frequency bands: $\delta$, $\theta$, $\alpha$, $\beta$, $\gamma$. Frequency domain features, such as power spectral density (PSD) [3], differential entropy (DE) [4, 5], differential asymmetry (DASM) [6] and rational asymmetry (RASM) [7], are subsequently extracted from each frequency band respectively.
@@
The spatial structure of the brain also contains rich emotional information. Emotional states may involve distributed circuits rather than only a single brain region [8]. Asymmetry between the left and right hemispheres can reflect changes in valence and arousal [9]. Recent studies have highlighted the importance of utilizing spatial domain information. Li et al. [10] introduced recurrent neural networks to learn the asymmetric differences between the left and right hemispheres. Li et al. [11] also utilized hierarchical neural networks to learn both regional and global information of spatial-temporal EEG features. Graph neural networks (GNNs) are emerging as a powerful tool for analyzing spatial information in EEG emotion recognition. Song et al. [12] dynamically learned relationships between EEG channels using a graph convolutional network (GCN). Zhong et al. [13] incorporated asymmetry of the hemispheres into the adjacency matrix to model graph structure. Li et al. [14] also utilized an multi-domain adaptive graph convolutional network (MD-AGCN) to learn relationships between channels. Ding et al. [15] incorporated lobe information as prior knowledge into the GNN. Jiang et al. [16] proposed an elastic graph transformer (EmoGT) to extract emotional information. Although these methods have achieved excellent performance in emotion recognition, they have a common issue: they all rely on GNNs based on neighborhood aggregation schemes which may pose potential risks such as over-smoothing [17, 18, 19], under-reaching [20], and over-squashing [21]. Additionally, these methods do not take the geometric and anatomical structure information of the brain into account.
@@
The main contributions of this paper lie in three aspects: (1) We propose a multi-view graph transformer (MVGT) based on spatial relations, fusing information from multiple perspectives including temporal, frequency, and spatial domains. (2) Our method, based on graph transformer, mitigates the potential risks of over-smoothing, under-reaching and over-squashing occurring in traditional GNNs. Additionally, it enhances the model’s expressive power in emotion recognition by introducing spatial encodings based on geometric and brain lobe information. (3) Extensive experiments conducted on public datasets SEED and SEED-IV show our model achieves superior performance over the baseline models in emotion classification tasks.
@@ II Related Work
In this section, we review the related work in terms of EEG-based emotion recognition and graph transformer.
@@ II-A EEG-based Emotion Recognition
Along with the development of deep learning, increasingly complex models with rich expressive abilities have emerged and have been extensively utilized in EEG signal analysis. Zheng et al. [5] employed a deep belief network to analyze important frequency domain components and effective channels based on the learned parameters. Song et al. [12] used a GCN method based on Chebyshev polynomials [24] to dynamically learn the representations of EEG signals. Zhong et al. [13] innovatively incorporated the inter-channel asymmetry of the hemispheres as prior knowledge into the adjacency matrix in 3D space. The reasonable combination of multi-domain information improves the accuracy in the emotion recognition tasks. Li et al. [14] proposed the MD-AGCN that integrates the temporal domain, frequency domain, and functional connectivity. Ding et al. [15], inspired by neuroscience research, combined intra-region convolution and inter-region convolution based on brain lobe information to learn cognitive patterns. Jiang et al. [16] utilized the advantages of GCN in the spatial domain and Transformer in the temporal domain to improve the accuracy of emotion classification.
@@ II-B Graph Transformer
The GNNs used in the methods above are based on neighborhood aggregation schemes. However, classical GNNs based on message passing (MPGNNs) may lead to over-smoothing [17, 18, 19], under-reaching [20], and may also fail to fit long-range signals due to over-squashing [21], which limit the expressive power of the model. Graph transformers (GTs) alleviate such effects as they have a global receptive field [25]. Nevertheless, without sufficiently expressive structural and positional encodings, GTs cannot capture effective graph structures [26]. Dwivedi et al. [27] utilized eigenvectors of graph Laplacian as position encodings in a fully connected graph transformer and integrated edge features into the attention mechanism. Building on this, SAN [28] used a full Laplacian spectrum to learn the positional encodings for each node. Graphormer [29, 30] employed node centrality and node distance metric to implement structural and relative positional encodings, achieving state-of-the-art performance on challenging graph datasets. For EEG emotion recognition, Li et al. [31] innovatively combined a masked autoencoder based on self-supervised learning with a CNN-Transformer hybrid structure, effectively improving classification accuracy. However, this method only used the sine-cosine positional encoding, limiting the Transformer’s ability to learn spatial information.
@@ III-A Graph Neural Network (GNN)
Let $G=\left(V,E\right)$ define a graph, where $V=\left\{v_{1},v_{2},\cdots,v_{n}\right\}$ represents the nodes in the graph, and $E=\left\{e_{1},e_{2},\cdots,e_{m}\right\}$ is the edges between the nodes. The representation of node $v_{i}$ is denoted as $x_{i}\in\mathbb{R}^{d}$. Most existing GNNs [32, 17, 33, 34, 35] adopt neighborhood aggregation schemes, iteratively aggregating representations of its first or higher-order neighbors, followed by using backpropagation (BP) to learn data-driven feature representations. We define the representation of node $v_{i}$ at the $l$-th iteration as $h_{i}^{l}$ and define $h_{i}^{0}=x_{i}$. The $l$-th iteration can be represented as:
$a_{i}^{l}$ $=\text{AGGREGATE}^{l}\left(\left\{\varphi_{\theta}(h_{j}^{l-1},e_{ji}):j\in\mathcal{N}(v_{i})\right\}\right),$ (1)
$h_{i}^{l}$ $=\text{UPDATE}^{l}\left(h_{i}^{l-1},a_{i}^{l}\right),$ (2)
where $\varphi_{\theta}$ represents a differentiable function used for feature transformation of node and edge information. The set $\mathcal{N}(v_{i})$ is the neighbors of $v_{i}$. The AGGREGATE function is used to aggregate the transformed representations using a differentiable, permutation invariant function (such as mean, sum, max, etc.). The goal of UPDATE function is to integrate the information from neighbors into the node representation. For graph classification, the READOUT operation is typically used to obtain a representation of the entire graph, which is then fed into a classifier to determine the graph label.
@@ III-B Graphormer
The Transformer [36] is undeniably one of the most popular deep neural network architectures today, driving significant advancements in natural language processing and computer vision. From the perspective of GNNs, Transformer can be interpreted as a GNN acting on a fully connected graph. Therefore, it is feasible to use the Transformer to address tasks on graph data. The ability to properly incorporate the structural information of graphs into the model is the key for leveraging its expressive power. Graphormer [29, 30] can go beyond classical MPGNNs in expressive power and achieves state-of-the-art performance on large molecular benchmarks. Graphormer incorporates centrality encoding into the graph data and integrates spatial encoding, edge encoding into the attention mechanism, which can be expressed as:
$\bm{A}_{ij}=\frac{\left(\bm{h}_{i}\bm{W}_{Q}\right)\left(\bm{h}_{j}\bm{W}_{K}\right)^{\top}}{\sqrt{d}}+b_{\phi(v_{i},v_{j})}+c_{ij},$ (3)
@@ IV METHODS
In this section, we introduce the methods employed in the EEG emotion recognition task. First, we elaborate on the embedding of temporal information. Second, leveraging the spatial geometry and physiological anatomy of the brain, we propose two novel and simple designs of encoding that enable the model to adaptively learn the inter-channel correlations. Finally, we present the detailed implementations of MVGT.
@@ IV-A Problem Definition
EEG signals can be represented as a two-dimensional matrix with respect to channels and time. Given that channels exhibit spatial structure, they can be structured into a fully connected graph $G=\left(V,E\right)$, where $V$ denotes the nodes, representing EEG channels, and $E$ denotes the edges, representing the connections between channels. The features of the nodes at time $t$ are denoted by $\bm{X}_{t}=\left[\bm{x}_{1}^{\top},\bm{x}_{2}^{\top},\cdots,\bm{x}_{n}^{\top}\right]^{\top}\in\mathbb{R}^{n\times d}$, where $n=|V|$ represents the number of nodes and $d$ represents the feature dimension.
@@ IV-B Temporal Embedding
EEG signals have high temporal resolution and contain rich temporal information. Because of the multi-electrode acquisition method, EEG signals can be regarded as multivariate time series. When processing time series, the embedding of temporal information are crucial. EmoGT treated the features of different channels at the same time points as tokens and employed an attention mechanism to extract temporal correlations between them. Due to the anisotropic volume conduction characteristics [37] in human brain tissues, there may be temporal delays between different channels, which in turn leads to time-unaligned events at a single moment thus causing performance degradation. MD-AGCN utilized the convolutional operation to extract temporal information along the time axis from continuous EEG segments, with the receptive field limited by the size of the convolution kernel. Inspired by [38], we broaden the receptive field by considering the entire time series as an embedded token rather than a single time point. First, following the methods of MD-AGCN and EmoGT, we use overlapping sliding windows of size $T$ to split EEG signals along the time axis and use these segments as samples, which are then fed into the attention module in the form of continuous segments. After processing with sliding windows, we obtain $\tilde{\bm{X}}=\left(\tilde{\bm{X}}_{1},\tilde{\bm{X}}_{2},\cdots,\tilde{\bm{X}}_{S}\right)$, where the $s$-th sample is $\tilde{\bm{X}}_{s}\in\mathbb{R}^{n\times Tf}$, $S$ is the number of continuous EEG segments, $n$ is the number of channels, and $f$ is the dimension of frequency domain features.
@@ IV-C Spatial Encoding
The special structure of the brain encompasses rich spatial information. Fully exploiting structural information is beneficial for the recognition and analysis of cognitive patterns in the brain. Therefore, to better recognize emotional patterns in emotion classification tasks, we employ two simple but effective methods of spatial encoding: brain region encoding and geometric structure encoding.
@@ IV-C1 Brain Region Encoding
Neuroscience research demonstrated that the activation of a specific brain region often leads to the concurrent activation of related brain regions responsible for the same high-level cognition [40]. In EEG emotion recognition, incorporating relevant neuroscience findings can typically enhance recognition accuracy. Zhong et al. [13] integrated the asymmetry of neural activity between the left and right hemispheres as prior knowledge into the adjacency matrix, effectively enhancing recognition accuracy. Li et al. [10] improved emotion pattern recognition performance by capturing the differences between the left and right hemispheres. Ding et al. [15] divided EEG channels into different regions and combined local intra-region convolution with global inter-region convolution, achieving good results on the DEAP [41] dataset. With reference to the three divisions of [15], we adopt four brain region divisions, which divide the EEG channels into different regions based on a prior knowledge, aiming to incorporate the brain region information into the model. The division schemes are as follows:
  • $\bullet$
    We divide the regions based on the anatomical structure of the brain and implement LOBE scheme.
  • $\bullet$
    To further investigate the expressive power of brain region encoding, we conduct a detailed division of brain lobes according to the 10-20 system based on electrode positions, employing the GENERAL scheme.
  • $\bullet$
    Asymmetric EEG activity in the frontal lobe can be utilized for discriminating valence changes [9]. The left frontal lobe exhibits a stronger correlation with joy and happiness, while the right frontal lobe is more strongly correlated with fear and sadness. Thus we further divide the frontal lobe region into two symmetrical regions to obtain the FRONTAL scheme.
  • $\bullet$
    According to the symmetry of brain structure [42], we make a finer division of the brain lobe regions, defining the HEMISPHERE scheme.
The four schemes mentioned above are shown in Fig. 2. In terms of specific implementation, we assign a brain region tag to each electrode, then map the tags into an embedding space using a learnable projection function, and simply add the embeddings to the node features. The encoding of node $i$ is represented as follows:
@@
where $\bm{W}_{\mathcal{X}}\in\mathbb{R}^{Tf\times d}$ is a learnable projection function, and $d$ represents the dimension of the embedding. Through the above encoding method, we integrate the brain’s anatomy information into the model.